# STGFormer: Spatio-Temporal GraphFormer for 3D Human Pose Estimation in Video

### v1

@@ Abstract
The current methods of video-based 3D human pose estimation have achieved significant progress; however, they continue to confront the significant challenge of depth ambiguity. To address this limitation, this paper presents the spatio-temporal GraphFormer framework for 3D human pose estimation in video, which integrates body structure graph-based representations with spatio-temporal information. Specifically, we develop a spatio-temporal criss-cross graph (STG) attention mechanism. This approach is designed to learn the long-range dependencies in data across both time and space, integrating graph information directly into the respective attention layers. Furthermore, we introduce the dual-path modulated hop-wise regular GCN (MHR-GCN) module, which utilizes modulation to optimize parameter usage and employs spatio-temporal hop-wise skip connections to acquire higher-order information. Additionally, this module processes temporal and spatial dimensions independently to learn their respective features while avoiding mutual influence. Finally, we demonstrate that our method achieves state-of-the-art performance in 3D human pose estimation on the Human3.6M and MPI-INF-3DHP datasets.
Index Terms:
3D human pose estimation, vision transformer, graph convolution
I Introduction
@@
3D human pose estimation predicts the coordinates of body keypoints in three dimensions from input images or videos. Compared to 2D pose estimation, it provides more comprehensive and accurate spatial information. In the field of computer vision, this technology has garnered widespread attention due to its potential applications in action recognition [1, 2], human-computer interaction [3], and autonomous driving [4, 5]. 3D human pose estimation can primarily be divided into single-stage methods and two-stage methods. Single-stage methods directly estimate the end-to-end 3D human pose from inputs; two-stage methods initially detect the coordinates of keypoints in images and then regress to a 3D human pose based on the detected 2D keypoints. Currently, research shows that the performance metrics of two-stage methods generally surpass those of single-stage methods, mainly due to the significant advancements in 2D pose estimation networks. For instance, ViTPose [6] utilizes a transformer as the backbone to extract features and employs a lightweight decoder for pose estimation. The proposed method has been demonstrated to be effective and flexible in 2D pose estimation. However, the second stage, which involves transitioning from 2D pose to 3D, remains unsatisfactory due to the ill-posed problem. This issue stems from the depth ambiguity inherent in estimating 3D coordinates from 2D, caused by a lack of depth prior information.
@@
To address the aforementioned issues, a series of methods have been proposed that are based on backbones such as CNNs [7, 8, 9, 10], GCNs [11, 12, 13, 14, 15], and transformers [16, 17, 18, 19, 20, 21, 22]. For example, Einfalt et al. [19] perform temporal upsampling within Transformer blocks to process time-sparse 2D pose sequences, thereby reducing the model’s size while still generating time-dense 3D pose estimates. StridedFormer [22] reduces sequence redundancy by substituting the fully connected layers in the feedforward network of the Vanilla Transformer Encoder (VTE) with strided convolutions, thereby progressively decreasing the sequence length to lower computational costs. Although the use of transformers capitalizes on the advantages of scalable model capacity and high parallelism, which show promising prospects in enhancing the performance of pose estimation, several challenges still persist.
@@
First, owing to the absence of structural priors in the transformer, its performance in 3D human pose estimation still trails behind the state-of-the-art methods based on GCNs. These graph-based methods effortlessly represent the human skeleton as a graph, where joints are regarded as nodes, and the bones between joints are considered as edges connecting these nodes. Currently, some improvements that integrate graphs with transformers have emerged [23, 24, 25]. For instance, STCFormer [25] introduces body structure graph information as inductive biases into the transformers. This strategy influences the versatility of the transformers, which are commonly regarded as structures with weak inductive biases, with significant implications for pre-training and the scalability of large models.
Second, spatio-temporal transformers excel in capturing long-range dependencies, whereas GCN architectures integrate information and inter-node relationships. Encoding graph information into the attention layer significantly enhances the model’s understanding of direct connections between nodes. Meanwhile, replacing MLP with GCN in transformer may amplifie the ability to aggregate information throughout the entire graph structure. Integrating these two approaches allows the model to process graph data with greater precision and efficiency, thus optimizing performance from different aspects. However, in the most of GCN architectures, the transformation matrices are shared among all nodes within the multi-layer convolutional filters, impeding the efficiency of information exchange among nodes. To address this issue, mechanisms such as weight unsharing [26] and weight modulation [27] have been introduced in the spatial dimension, yet these solutions have not been effectively extended to the temporal dimension.
Third, higher-order dependencies are often neglected in the processing of graph neural networks. Although some studies address these dependencies in the spatial dimension by extending the first-order neighbors to high-order neighbors [28, 29], they remain unresolved in the temporal dimension. This issue of temporal higher-order dependencies originates from the system’s dynamic characteristics, whereby the current state is influenced not only by the immediately preceding state but also by earlier states.
To overcome the limitations discussed above, this paper presents a framework that integrates spatio-temporal transformer with graph encoding and dual-path modulated hop-wise regular GCN for 3D human pose estimation. This framework not only captures spatio-temporal long-range dependencies but also elucidates the relationships between spatio-temporal nodes and their structures. We summarize the main contributions of this work as follows.
(1) We propose a spatio-temporal GraphFormer for 3D human pose estimation from monocular video using spatio-temporal criss-cross graph attention and dual-path modulated hop-wise regular GCN.
(2) We introduce a spatio-temporal criss-cross graph attention mechanism which encodes temporal and spatial graph structural relationships into attention layers.
(3) We design a dual-path modulated hop-wise regular GCN module and integrate it into our GraphFormer architecture.
(4) We demonstrate extensive experiments conducted on Human3.6M and MPI-INF-3DHP datasets and ablation studies that our proposed model achieves state-of-the-art performance in comparison with strong baselines.
II Related Work
@@
In recent years, graph convolutional networks and transformers have made significant progress in 3D human pose estimation. As previously mentioned, we implement 3D human pose estimation from image sequences or video inputs through a two-stage approach.
II-A GCN-based methods.
@@
The skeletal structure of body can be naturally represented as a graph, hence there has been extensive research in image-based 3D human pose estimation using GCNs, which has yielded significant results [13, 15, 26, 27, 30, 31]. To more effectively address the ill-posed problem of lifting 2D joint keypoints into 3D space, additional temporal information is extracted from sequential frames combined with spatial information related to body structure. For instance, the Global-local Adaptive Graph Convolutional Network (GLA-GCN) [11] models the spatio-temporal structure using a graph representation and traces local joint features through individually connected layers. Zeng et al. [32] developed a hop-aware hierarchical channel-squeezing fusion mechanism to efficiently extract relevant information from neighboring nodes in GCNs. Hu et al. [33] represent the human skeleton as a directed graph to effectively reflect the hierarchical relationships among the nodes. Although graph-based methods effectively utilize the body’s structural information, they encounter a significant limitation in terms of scalability. Adding more network layers does not enhance model performance, and even widely adopted residual connections yield minimal impact.
II-B Transformer-based methods.
In addition to GCN, transformer also used to model spatio-temporal correlations for video-based 3D human pose estimation [17, 18, 19, 20, 21, 22, 34, 35]. Specifically, MixSynthFormer [17] combines synthesized spatial and temporal attention to effectively integrate inter-joint and inter-frame significance. HSTFormer [18] captures multi-level joint spatial-temporal correlations, ranging from local to global, through four transformer encoders and a fusion module. Li et al. [22] utilized a vanilla transformer encoder to model long-range dependencies in 2D pose sequences, while a strided transformer encoder aggregates this long-range information into a single-vector representation, employing a hierarchical global and local approach.
@@
The Transformer has high scalability and is good at capturing long-distance dependencies, but it lacks structural prior information. As a result, several methods have been developed that integrate graph and transformer. The Spatio-Temporal Criss-cross Attention (STC) [25] learns spatio-temporal features and incorporates the body’s graph structure information as a biased input in the proposed framework. Zhai et al. [24] proposed the Hop-wise GraphFormer, which enhances GCN-based pose estimation networks by grouping joints according to k-hop neighborhoods and capturing potential joint correlations across various joint synergies. Zhao et al. [23] designed a novel model, GraFormer, that combines graph convolution and transformers to exploit relationships among graph-structured 2D joints.
@@
Our work also belongs to the realm of 3D pose lifting that employs a fusion of transformer and graph. Unlike other approaches, our model integrates graph information directly into the attention layers, thereby significantly improving adaptability. Furthermore, it incorporates spatio-temporal hop-wise modulated GCNs within the transformer framework, effectively capturing high-order dependencies in both time and space. This approach complements the long-range dependencies typical of transformers, thus substantially enhancing the model’s capacity to represent complex data structures.
III Method
@@ III-A Preliminary
First, let us clarify the problem that needs to be solved in 3D human estimation based on deep learning. The training set for 3D human pose estimation contains paired data $\mathcal{D}={(\mathbf{x}_{i},\mathbf{y}_{i})},{i=1,\cdots,N}$, including 2D joint positions $\mathbf{X}=(\mathbf{x}_{1},\cdots,\mathbf{x}_{N})^{\mathsf{T}}\in\mathbb{R}^{N\times 2}$ and their associated 3D joint positions $\mathbf{Y}=(\mathbf{y}_{1},\cdots,\mathbf{y}_{N})^{\mathsf{T}}\in\mathbb{R}^{N\times 3}$. The goal of human pose lifting is to minimize the loss function $l(f(\mathbf{x}_{i}),\mathbf{y}_{i})$ of a regression model $f:\mathbf{X}\rightarrow\mathbf{Y}$ by learning the parameters.
@@
Then, we provide some preliminary knowledge regarding human pose estimation using graph learning methods. Consider a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}={1,\cdots,N}$ denotes the set of $N$ nodes and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ represents the set of edges. In the context of 3D human pose estimation, nodes correspond to human joints, and edges represent the connections between these joints. The adjacency matrix $\mathbf{A}=(a_{ij})$, with dimensions $N\times N$, contains elements $a_{ij}$ that represent the weights between node $i$ and node $j$. These weights may be expressed as binary or real numbers. In our approach, we utilize a binary representation, where a value of 1 indicates a connection $i\sim j$ between joints, while a value of 0 signifies the absence of such a connection.
The feature processing can transform low-dimensional joints into high-dimensional features with a dimension of $F$. Each joint can be represented by an $N\times F$ feature matrix $\mathbf{H}=(\mathbf{h}_{1},\cdots,\mathbf{h}_{N})^{\mathsf{T}}$, where each $\mathbf{h}_{i}$ is an $F$-dimensional feature vector corresponding to node $i$. For the joint inputs across consecutive frames $T$, they can be represented by a $T\times N\times F$ feature matrix $\mathbf{H}\in\mathbb{R}^{T\times N\times F}$. In the vanilla GCN [36], a graph convolutional layer transforms and aggregates the input features according to the equation below:
$\mathbf{H}^{\prime}=\sigma(\mathbf{WH}\tilde{\mathbf{A}})$ (1)
where the $i$-th row of the output feature matrix can be expressed as follows:
$\mathbf{h}_{i}^{\prime}=\sigma(\sum_{j\in\mathcal{N}_{i}}{\mathbf{Wh}_{j}}\tilde{a}_{ij})$ (2)
But the shared feature transformation weight $\mathbf{W}$ in the vanilla GCN restricts its ability to learn diverse relational patterns from different body joint nodes. Liu et al. [26] address this limitation by utilizing a unique weight matrix $\mathbf{W}_{j}\in\mathbb{R}^{F^{\prime}\times F}$ for each node $j$, as represented by the following equation:
$\mathbf{h}_{i}^{\prime}=\sigma(\sum_{j\in\mathcal{N}_{i}}{\mathbf{W}_{j}}\mathbf{h}_{j}\tilde{a}_{ij})$ (3)
To reduce the model size, weight modulation [27] in a shared feature transformation for different nodes is introduced:
$\mathbf{H}^{\prime}=\sigma((\mathbf{M}\odot(\mathbf{WH}))\tilde{\mathbf{A}})$ (4)
$\mathbf{h}_{i}^{\prime}=\sigma(\sum_{j\in\mathcal{N}_{i}}{(}\mathbf{m}_{j}\odot\mathbf{W})\mathbf{h}_{j}\tilde{a}_{ij})$ (5)
And then, to enhance feature reuse in GCN and prevent falling into local optima, RSNet [13] implements residual regular connections across different layers:
$\mathbf{H}^{\prime}=\sigma\left(\tilde{\mathbf{A}}((\mathbf{H}\mathbf{W})\odot\mathbf{M})+\mathbf{X}\tilde{\mathbf{W}}\right)$ (6)
$\mathbf{h}_{i}^{\prime}=\sigma\left(\sum_{j\in\mathcal{N}_{i}}{\tilde{a}_{ij}}((\mathbf{h}_{j}\mathbf{W})\odot\mathbf{m}_{j})+\mathbf{x}_{i}\widetilde{\mathbf{W}}\right)$ (7)
@@
Finally, it is essential to introduce briefly the transformer, which serves as the backbone of our framework. The transformer architecture is composed of a set of transformer layers, each consisting of two components: a self-attention module and a position-wise feed-forward network. Let $\mathbf{H}\in\mathbb{R}^{T\times N\times F}$ denote the input of the self-attention module where $F$ is the hidden dimension and $\mathbf{h}_{t,i}\in\mathbb{R}^{1\times 1\times F}$ is the hidden representation at joint $i$ in the frame $t$. The input $\mathbf{H}$ is projected by three matrices $\mathbf{W}_{Q}\in\mathbb{R}^{d\times d_{K}}$, $\mathbf{W}_{K}\in\mathbb{R}^{d\times d_{K}}$ and $\mathbf{W}_{V}\in\mathbb{R}^{d\times d_{V}}$ to the corresponding representations $Q$, $K$, $V$. The self-attention is then calculated as:
@@
where $\mathbf{A}$ is a matrix capturing the similarity between queries and keys.
III-B Overall Architecture
Figure 1(a) provides an overview of the proposed framework, which mainly comprises three parts: a joint-based embedding, spatio-temporal criss-cross graph transformer blocks, and a regression head. The joint-based embedding projects the input 2D coordinates of each joint into the feature space. STGFormer blocks aggregate the spatio-temporal context and graph information, updating the representation of each joint. The 3D coordinates are estimated from the learnt features by a regression head.

### v2

@@ Abstract
The current methods of video-based 3D human pose estimation have achieved significant progress. However, they still face pressing challenges, such as the underutilization of spatiotemporal body-structure features in transformers and the inadequate granularity of spatiotemporal interaction modeling in graph convolutional networks, which leads to pervasive depth ambiguity in monocular 3D human pose estimation. To address these limitations, this paper presents the Spatio-Temporal GraphFormer framework (STGFormer) for 3D human pose estimation in videos. First, we introduce a Spatio-Temporal criss-cross Graph (STG) attention mechanism, designed to more effectively leverage the inherent graph priors of the human body within continuous sequence distributions while capturing spatiotemporal long-range dependencies. Next, we present a dual-path Modulated Hop-wise Regular GCN (MHR-GCN) to independently process temporal and spatial dimensions in parallel, preserving features rich in temporal dynamics and the original or high-dimensional representations of spatial structures. Furthermore, the module leverages modulation to optimize parameter efficiency and incorporates spatiotemporal hop-wise skip connections to capture higher-order information. Finally, we demonstrate that our method achieves state-of-the-art performance on the Human3.6M and MPI-INF-3DHP datasets.
keywords:
[…]
1 Introduction
@@
3D human pose estimation predicts the coordinates of body keypoints in three dimensions from input images or videos. Compared to 2D pose estimation, it provides more comprehensive and accurate information about spatial relationships. In the field of computer vision, this technology has garnered widespread attention due to its potential applications in action recognition [1, 2], human-computer interaction [3], and autonomous driving [4, 5]. Monocular 3D human pose estimation can primarily be divided into single-stage methods and two-stage methods. Single-stage methods directly estimate the end-to-end 3D human pose from inputs; two-stage methods initially detect the coordinates of keypoints and then regress to a 3D human pose based on the detected 2D keypoints. Currently, research indicates that the performance metrics of two-stage methods generally outperform those of single-stage methods, mainly due to the significant advancements in 2D pose estimation networks. For instance, ViTPose [6] utilizes a transformer as the backbone to extract features and employs a lightweight decoder for pose estimation. The proposed method has been demonstrated to be effective and flexible in 2D pose estimation. However, the second stage, which involves transitioning from 2D pose to 3D, remains unsatisfactory due to the ill-posed problem. This issue stems from the depth ambiguity inherent in estimating 3D coordinates from 2D, caused by a lack of depth prior information.
@@
To address the aforementioned issues, a series of two-stage approaches for monocular 3D human pose estimation have been proposed that are based on backbones such as Convolutional Neural Networks (CNN) [7, 8, 9, 10], Graph Convolutional Networks (GCN) [11, 12, 13, 14, 15], and transformers [16, 17, 18, 19, 20, 21, 22]. For example, Einfalt et al. [19] perform temporal upsampling within transformer blocks to process time-sparse 2D pose sequences, thereby reducing the model’s size while still generating time-dense 3D pose estimates. StridedFormer [22] reduces sequence redundancy by substituting the fully connected layers in the feedforward network of the Vanilla Transformer Encoder (VTE) with strided convolutions, thereby progressively decreasing the sequence length to lower computational costs. While transformers take advantage of the scalable capacity and high parallelism, offering promising improvements for pose estimation, they still face several challenges.
@@
First, owing to the absence of structural priors in the transformer, its performance in 3D human pose estimation still trails behind the state-of-the-art methods based on GCNs. These graph-based methods effortlessly represent the human skeleton as a graph, where joints are regarded as nodes, and the bones between joints are considered as edges connecting these nodes. Currently, some improvements that integrate graphs with transformers have emerged [23, 24, 25]. For instance, STCFormer [25] introduces a criss-cross attention strategy that, while mining the correlations across frame-level features, explores the relationships between joints across different frames. However, this method incorporates body structure graphs as inductive biases into the transformers. This approach affects the versatility of transformers, which are typically regarded as architectures with weak inductive biases, and has significant implications for pre-training and the scalability of large models.
To address the existing limitations, we propose a spatio-temporal criss-cross graph attention mechanism as part of our STGFormer architecture. This mechanism seamlessly integrates spatiotemporal graph structures into the attention layers, allowing the model to leverage prior knowledge about the human skeleton more effectively. By doing so, our approach enhances the ability to capture long-range dependencies across both spatial and temporal dimensions, ultimately improving performance in 3D human pose estimation.
Second, replacing Multi-Layer Perceptron (MLP) with GCN in transformer may enhance the ability to aggregate information throughout the entire graph structure. Integrating these two approaches allows the model to process graph data with greater precision and efficiency, thus optimizing performance from different aspects. However, in GCN computations, the computational cost grows exponentially with an increasing number of frames, rendering the approach impractical.
In this paper, we decouple temporal feature learning and spatial feature learning into a dual-path parallel operation within our MHR-GCN, which reduces the computational complexity from $\mathcal{O}\left(S\cdot T+(S-1)(T-1)\right)$ to $\mathcal{O}\left(S+(S-1)\right)+\mathcal{O}\left(T+(T-1)\right)$, where $S$ and $T$ denote the number of joints and frames, respectively. Moreover, the parallel dual-path module circumvents the sequential processing in serial architectures, which can lead to the premature compression of certain features. This approach effectively preserves both temporally dynamic features and the original or high-dimensional representations of spatial structures.
Third, in image-based 3D human pose estimation, mechanisms such as weight unsharing [26] and weight modulation [27] have been adopted to improve the efficiency of information exchange among nodes by preventing the use of uniformly shared transformation matrices. Meanwhile, higher-order dependencies have received increasing attention, with several studies addressing these dependencies by extending connections from first-order to higher-order neighbors [28, 29]. Temporal higher-order dependencies originate from the system’s dynamic characteristics, whereby the current state is influenced not only by the immediately preceding state but also by earlier states. However, the adoption of these methods in the temporal dimension has been neglected in video-based 3D human pose estimation.
To address this gap, we propose a novel spatiotemporal adjacency modulation mechanism within our MHR-GCN architecture, extending the original spatial modulation to the temporal dimension. This module further extends the original spatial modulation to the temporal dimension, preventing reliance on static or uniformly shared transformation matrices and thereby improving the flow of information across the network. In addition, we incorporate hop-wise spatiotemporal connections, which extend beyond immediate neighbors to capture higher-order dependencies across both space and time. By combining adjacency modulation with hop-wise connections, our MHR-GCN robustly models the dynamic characteristics of human motion and more effectively utilizes temporal higher-order dependencies in video-based 3D human pose estimation.
Our main contributions are summarized as follows:
(1) We propose a spatiotemporal criss-cross graph attention mechanism within our STGFormer architecture, integrating spatiotemporal graph structures into attention layers to more effectively leverage prior knowledge when capturing long-range dependencies.
(2) We introduce a dual-path modulated hop-wise regular GCN module and incorporate it into STGFormer, extracting first-order and higher-order spatiotemporal graph information from the features.
(3) We conduct extensive experiments on Human3.6M [30] and MPI-INF-3DHP [31] datasets for a comprehensive evaluation, showing that our proposed model performs well against the state-of-the-art methods.
2 Related Work
@@
In recent years, both graph convolutional networks and transformers have made significant strides in 3D human pose estimation. As previously mentioned, we apply a two-stage approach to perform 3D human pose estimation from monocular image sequences or video inputs.
2.1 GCN-based methods.
@@
The skeletal structure of the human body can be naturally represented as a graph. As a result, extensive research has been conducted on image-based 3D human pose estimation using GCNs, yielding significant results [13, 15, 26, 27, 32, 33]. To more effectively address the ill-posed problem of lifting 2D joint keypoints into 3D space, additional temporal information is extracted from sequential frames combined with spatial information related to body structure. For instance, the Global-Local Adaptive Graph Convolutional Network (GLA-GCN) [11] models the spatio-temporal structure using a graph representation and traces local joint features through independently connected layers. Zeng et al. [34] developed a hop-aware hierarchical channel-squeezing fusion mechanism to efficiently extract relevant information from neighboring nodes in GCNs. Hu et al. [35] represent the human skeleton as a directed graph to effectively reflect the hierarchical relationships among the nodes. Although graph-based methods effectively leverage the body’s structural information, they face a significant limitation in scalability. Adding more network layers does not improve model performance, and even commonly used residual connections have minimal impact.
2.2 Transformer-based methods.
@@
In addition to GCN, transformers are also used to model spatio-temporal correlations for video-based 3D human pose estimation [17, 18, 19, 20, 21, 22, 36, 37]. Specifically, MixSynthFormer [17] combines synthesized spatial and temporal attention to effectively integrate inter-joint and inter-frame significance. HSTFormer [18] captures multi-level joint spatial-temporal correlations, ranging from local to global, through four transformer encoders and a fusion module. Li et al. [22] utilized a vanilla transformer encoder to model long-range dependencies in 2D pose sequences, while a strided transformer encoder aggregates this long-range information into a single-vector representation, employing a hierarchical global and local approach. Transformers exhibit high scalability and are proficient at capturing long-range dependencies; however, they lack structural prior information. As a result, several methods have been developed that integrate graph and transformer. The Spatio-Temporal Criss-cross (STC) attention [25] learns spatio-temporal features and incorporates the body’s graph structure information as a biased input in the proposed framework. Zhai et al. [24] proposed the Hop-wise GraphFormer, which enhances GCN-based pose estimation networks by grouping joints according to k-hop neighborhoods and capturing potential joint correlations across various joint synergies. Zhao et al. [23] designed a novel model, GraFormer, that combines graph convolution and transformers to exploit relationships among graph-structured 2D joints.
@@
Our work also falls within the realm of 3D pose lifting, which employs a fusion of transformers and graphs. Unlike other approaches, our model directly integrates graph information into the attention layers, thereby significantly enhancing adaptability. Furthermore, it incorporates spatio-temporal hop-wise modulated GCNs within the transformer framework, effectively capturing high-order dependencies in both time and space. This approach complements the long-range dependencies typical of transformers, thus substantially enhancing the model’s capacity to represent complex data structures.
3 Method
@@ 3.1 Preliminary
First, let us clarify the problem that needs to be solved in two-stage 3D human estimation based on deep learning. The training set for 3D human pose estimation contains paired data of $N$ joints $\mathcal{D}={(\mathbf{x}_{i},\mathbf{y}_{i})},{i=1,\cdots,N}$, including 2D joint positions $\mathbf{X}=(\mathbf{x}_{1},\cdots,\mathbf{x}_{N})^{\mathsf{T}}\in\mathbb{R}^{N\times 2}$ and their corresponding 3D joint positions $\mathbf{Y}=(\mathbf{y}_{1},\cdots,\mathbf{y}_{N})^{\mathsf{T}}\in\mathbb{R}^{N\times 3}$. The goal of human pose lifting is to minimize the loss function $l(f(\mathbf{x}_{i}),\mathbf{y}_{i})$ of a regression model $f:\mathbf{X}\rightarrow\mathbf{Y}$ by learning the parameters.
@@
Then, we provide some preliminary knowledge on human pose estimation using graph learning methods. Consider a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}={1,\cdots,N}$ denotes the set of $N$ nodes and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ represents the set of edges. In the context of 3D human pose estimation, nodes correspond to human joints, and edges represent the connections between these joints. The adjacency matrix $\mathbf{A}=(a_{ij})$, with dimensions $N\times N$, contains elements $a_{ij}$ that represent the weights between node $i$ and $j$. These weights may be expressed as binary or real numbers. In our approach, we utilize a binary representation, where a value of $1$ indicates a connection $i\sim j$ between joints, while a value of 00 signifies the absence of such a connection. The feature processing step transforms low-dimensional joints into high-dimensional features with a dimension of $F$. Each body can be represented by an $N\times F$ feature matrix $\mathbf{H}=(\mathbf{h}_{1},\cdots,\mathbf{h}_{N})^{\mathsf{T}}$, where each $\mathbf{h}_{i}$ is an $F$-dimensional feature vector corresponding to node $i$. For the inputs across consecutive frames $T$, they can be represented by a $T\times N\times F$ feature matrix $\mathbf{H}\in\mathbb{R}^{T\times N\times F}$.
@@
Finally, it is essential to briefly introduce the transformer, which serves as the backbone of our framework. The transformer architecture consists of a set of transformer layers, each consisting of two components: a self-attention module and a position-wise feed-forward network. Let $\mathbf{H}\in\mathbb{R}^{T\times N\times F}$ denote the input to the self-attention module where $F$ is the hidden dimension and $\mathbf{h}_{t,i}\in\mathbb{R}^{1\times 1\times F}$ is the hidden representation at joint $i$ in the frame $t$. The input $\mathbf{H}$ is projected by three matrices $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$ and $\mathbf{W}_{V}$ to the corresponding representations query $\mathbf{Q}$, key $\mathbf{K}$, and value $\mathbf{V}$. The self-attention is then computed as:
@@
where $\mathbf{A}$ is a matrix capturing the similarity between $\mathbf{Q}$ and $\mathbf{K}$, $\sqrt{d_{K}}$ represents the feature dimension of the query matrix.
3.2 Overall Architecture
Figure 1(a) provides an overview of the proposed framework, which mainly comprises three parts: a joint-based embedding, spatio-temporal criss-cross graph transformer blocks, and a regression head. The joint-based embedding projects the input 2D coordinates of each joint into a higher-dimensional feature space. The STGFormer blocks integrate the STG attention and the dual-path MHR-GCN module to capture spatiotemporal and graph features, thereby updating the representation of each joint. The 3D coordinates are estimated from the learned features by a regression head.